Dividing Partnership Income

Morrison and Greene have decided to form a partnership. They have agreed that Morrison is to invest $201,000 and that Greene is to invest $67,000. Morrison is to devote one-half time to the business, and Greene is to devote full time. The following plans for the division of income are being considered:

  1. Equal division.
  2. In the ratio of original investments.
  3. In the ratio of time devoted to the business.
  4. Interest of 5% on original investments and the remainder equally
  5. Interest of 5% on original investments, salary allowances of $60,000 to Morrison and $75,000 to Greene, and the remainder equally
  6. Plan (e), except that Greene is also to be allowed a bonus equal to 20% of the amount by which net income exceeds the total salary allowances

Required:

For each plan, determine the division of the net income under each of the following assumptions: (1) net income of $163,000 and (2) net income of $260,000. Round answers to the nearest whole dollar.

(1) (2)
$163,000 $260,000
Plan Morrison Greene Morrison Greene
a. $81,500 $81,500 $130,000 $130,000
b. $122,250 $40,750 $195,000 $65,000
c. $54,333 $108,667 $86,667 $173,333
d. $84,850 $78,150 $133,350 $126,650
e. $77,350 $85,650 $125,850 $134150
f. $ $ $ $
Answer #1

Answer :

Morrison Greene Morrison Greene
(f) 201000*5%+60,000+7020 77070 67000*5%75,000+560+7020 85930 Net income - total salary = 163000-60000-75000 = 28000,Bonus to green = 28000*2% = 560,remainder 163000-10050-3350-60000-75000-560= 14040,divided by 2 = 14040/2 = 7020 201000*5%+60,000+54550 124600 67000*5%+75,000+54550+2500 135400 Net income - total salary = 260000-60000-75000 = 125000,Bonus to green = 125000*2%=2500,remainder 260000-10050-3350-60000-75000-2500=109100, divided by 2 = 109100/2 = 54550

Please upvote Thank You !!!